The sum of the ages of Demetrius and Guadalupe is 64 years. 8 years ago Demetrius age was 2 times guadalupes age. How old is Demetrius now?
Found 2 solutions by josgarithmetic, greenestamps:
Answer by josgarithmetic(39797)

The site has so many of these. Students should be able to find a gigantic number of them already solved.

Answer by greenestamps(13334)



Here is one of many ways to set the problem up using formal algebra.

d = Demetrius' age
g = Guadalupe's age
d-8 = D's age 8 years ago
g-8 = G's age 8 years ago

(1) d+g = 64 [the sum of their ages is 64]
(2) d-8 = 2(g-8) [Demetrius' age 8 years ago was twice Guadalupe's age then]

Solve the pair of equation by your favorite method....

Without formal algebra....

The sum of their ages now is 64; the sum of their ages 8 years ago was 64-2(8) = 48.

When the sum of their ages was 48, Demetrius was twice as old as Guadalupe; that means their ages were 32 and 16.

That was 8 years ago, so now Demetrius is 40 years old.